Help please I don't no how to do these

The curved section of a horizontal highway is a ci
Answer #1

8) Friction force = centripetal force

umg = mv^2/r =========> v = sqrt(urg) = sqrt(0.4*760*9.8) = 54.58 m/s

9) the normal force from the road

10) Centripetal acceleration around a circle is a = v^2/R.
In the case of the loop, the acceleration must at least equal the acceleration due to gravity - at the top of the loop the centripetal acceleration is pointing up, gravity is pointing dowm, and the if gravity is greater then the net acceleration is down and away from the seat:
g = v^2/R
v = sqrt(gR) = sqrt(9.81 m/s^2 * 15m) = 12.13 m/s

So at v= 12.13 m/s the rider hits "freefall", and at greater speeds actually still feels pressed into their seat.